A dealer altered his discount from 25 to 20 percent . By what percent was the selling price altered.
Answer by Theo(13342)

this can get tricky, but if you keep track of what your are doing, you should be able to arrive at the correct answer.

assume the original retail price is 1 dollar.
the old selling price is 1 - .25 * 1 = .75.
the new selling price is 1 - .20 * 1 = .80.
the difference between the new selling price and the old selling price is .80 minus .75 = .05.
the ratio of the difference from the old selling price is .05/.75 = .06666666667.
the percent of the difference is 100 * the ratio of the difference which is equal to 6.66666667%.
that's equivalent to 6 and 2/3 percent.
that's the percent that the old selling price was altered to get to the new selling price.

let's say the original retail price was 200 dollars.
25% discount on 200 dollars is equal to 50 dollars.
the old selling price is 150 dollars.
20% discount on 200 dollars is equal to 40 dollars.
the new selling price is 160 dollars.
the difference between the new selling price and the old selling price is 10 dollars.
the percent difference from the old selling price is 10/150 = 1/15 = .0666666667 multiplied by 100 = 6.666666667 percent which is equal to 6 and 2/3 percent.